# IEEEabrv.bib

### v1

@@ IV-B Object Trajectory Diffusion
Our method decouples observation-action pairs in visuomotor policy learning by employing object pose trajectory as an intermediate representation. We model $p(T_{t:t+h}|T_{t})$ to infer a sequence of future object pose trajectory targeted for task accomplishment $T_{t:t+h}\in SE(3)$, from the current object pose $T_{t}\in SE(3)$.
@@
$A_{t}^{k-1}=\alpha\left(A_{t}^{k}-\gamma\epsilon_{\theta}(O_{t},A_{t}^{k},k)+\mathcal{N}(0,\sigma^{2}I)\right),$ (1)
The process starts with ${A}_{t}^{K}$ sampled from Gaussian noise and is iterated $K$ times. Here, $(\alpha,\gamma,\sigma)$ are the parameters of the denoising scheduler. This process is also called reverse process and detailed in previous work [ho2020denoising]. The score function $\epsilon_{\theta}$ is trained by the loss function
@@
In our framework, we model $p(T_{t+1}\mid T_{t})$ as the primary component for policy learning. Rather than predicting future actions, we synthesize the object pose trajectory for task progression. Additionally, our diffusion process is conditioned on the historical pose until now rather than raw sensory input. To adapt this method to our framework, we replace the observation input $O_{t}$ with the current and past object pose $T_{t}$. Furthermore, to obtain the object pose trajectory, $A_{t}$ is replaced by the transformation $T_{t+1}$. Accordingly, we modified Eq. (1) to:
$T_{t}^{k-1}=\alpha\left(T_{t}^{k}-\gamma\epsilon_{\theta}(T_{t},T_{t+1}^{k},k)+\mathcal{N}(0,\sigma^{2}I)\right),$ (3)
@@
Multi-task Language Conditioning. Thanks to the flexibility of diffusion models, we can effortlessly adapt to single-model multi-task settings by conditioning on other multi-modal data, such as language. Specifically, task descriptions are fed into a pretrained CLIP [radford2021learning] model to obtain sentence embeddings, following previous work [shridhar2023perceiver, goyal2023rvt, goyal2024rvt]. These feature embeddings are then concatenated with each object pose as input to the model.
We use DDIM [song2020denoising] as the noise scheduler. The scheduling is set to 100 timesteps during training and 10 timesteps during inference. The training epoch is 3000, and the batch size is 128 for all the baselines.
@@ IV-C Closed-loop Policy Execution
To derive the closed-loop manipulation policy $\pi$ from our object-centric representation, we need two key connections: one from observation to object pose, $O_{t}\mapsto T^{obj}_{cam}$, and another from pose trajectory to action plan, $T_{t:t+h}\mapsto A_{t:t+h}$. For the first mapping, we utilize off-the-shelf object pose tracking. For the second mapping, the action plan is derived from the synthesized object trajectory. The complete framework consists of three main components: Object Pose Tracking, which provides real-time object poses at each step; Trajectory synthesizing, which takes these poses as input and outputs the future object trajectory for task progression; and finally, Action plan generation, which derives the action plan from the synthesized object trajectory.
Object Pose Tracking. We utilize FoundationPose [wen2024foundationpose], a foundation model for 6D object pose estimation and tracking, which can be applied to novel objects in zero-shot. It tracks all the task-relevant SE(3) object poses throughout the policy rollout to provide the latest feedback to the controller. Besides RGBD observations, the tracker requires a 3D object model and a 2D detection for the first frame. The 3D object mesh model can be easily created using consumer-grade devices. In practice, we use an iPhone with LiDAR to scan the object into a textured mesh. For 2D object detection, we employ YOLO-World [cheng2024yolo], an open-vocabulary object detection model. It takes as input the RGB image of the scene and task-relevant object names as text prompts and outputs their 2D detections. We leverage ChatGPT to extract task-relevant object names from the task description automatically. Similar to the demonstration data processing, we convert the estimated source object pose into the target object’s coordinate frame before feeding it into the trajectory diffusion model.
@@
Trajectory synthesizing. During testing, we employ the trained diffusion model (Sec. IV-B) to generate future object trajectories targeted for task accomplishment. At each step, the model uses the previous (if the observation horizon is greater than 1) and current object poses as inputs to predict future object pose trajectories. We aim to achieve long-horizon planning while remaining reactive to unexpected dynamic uncertainties, such as in-hand object slipping or moving target objects. To this end, we integrate the object pose trajectories synthesized by the diffusion model with receding horizon control [mayne1988receding] to ensure robust action execution. Specifically, the model forecasts future object pose trajectories for $N$ steps, from which we select $K$ steps to execute actions. By adjusting $K$ and alternating between future trajectory prediction and trajectory following, we achieve closed-loop control based on updated observations from the object pose tracker.
@@
Action plan generation. The synthesized trajectory indicates the desired object pose to complete the task. To ensure the object achieves this pose, we need to convert the pose trajectory into an executable action plan. Specifically, we infer executable actions from the synthesized trajectory. Given the synthesized trajectory, we transform the future object poses into end effector subgoals using the transformation $T^{EE}_{obj}\in SE(3)$ from the object pose to the end effector pose. Given the extrinsic camera calibration, we can derive the transformation $T^{EE}_{obj}=T^{cam}_{EE}\cdot T^{obj}_{cam}\in SE(3)$ from the object pose to the end effector pose. Consequently, the action can be derived from the predicted object pose by $A_{t:t+h}=T^{EE}_{obj}\cdot T_{t:t+h}$. The action plan is constructed by following these end-effector subgoals. The policy continues to run until the episode ends and changes the gripper state, as signaled by the end-state prediction.
@@ V Experiments
Our experiments aim to address the following research questions: 1) Is our pose trajectory representation effective for capturing task-related 3D information? 2) Can our object-centric approach learn from action-less data and bridge the embodiment and environmental gap between training and testing? 3) Can our object-centric approach generate trajectories that comply with task-specific motion constraints? 4) Can we democratize the demonstration data collection with a minimal hardware setup?
@@ V-A Experimental Setup
Simulation setup. We evaluate baselines on RLBench [james2020rlbench], a standard multi-task manipulation benchmark adopted by previous studies [shridhar2023perceiver, goyal2023rvt, goyal2024rvt, ke20243d]. The task and the robot are simulated using CoppelaSim [rohmer2013v]. A Franka Panda robot equipped with a parallel jaw gripper is utilized for all tasks. Each task is accompanied by a language description and includes 2 to 60 variations, where the task-relevant objects are in different initial/goal configurations or appearances. We chose 13 manipulation tasks, excluding non-prehensile tasks and tasks with articulated objects, as they are beyond the scope of this work. Following [shridhar2023perceiver, goyal2023rvt], all methods are trained on the same set of 100 demonstrations and tested on 25 unseen task configurations.
Real-world setup. To validate learning from videos in unstructured settings without specific environmental or hardware requirements, we recorded demonstration videos in everyday environments using consumer-grade devices. Each task demonstration was performed by a human without robotics hardware, and filmed in various locations such as the kitchen, living room, or office, ensuring no overlap with the robot’s testing environment. This results in diverse visual backgrounds, lighting conditions, and camera setups that are different from those used during robot testing. The videos were captured with an iPhone equipped with RGB and LiDAR sensors, and all manipulations were performed by human hands. For each task, 8 demonstrations were collected for training, and 10 trials were conducted for each method during evaluation. All real-world experiments were conducted using the same hardware setup: a robot equipped with a Kinova Gen3 7 DoF arm and a forward-facing Intel Realsense 415 RGBD camera mounted at the left shoulder of the robot, which was extrinsically calibrated.
@@ V-B Simulation Results
Implementation details. To collect demonstration trajectories, we use the opening or closing of the gripper as signals to start or terminate a trajectory corresponding to an episode. For multi-stage tasks, in addition to language embedding, the object trajectory diffusion model is additionally conditioned on a one-hot encoding of the task stage. The next task stage is automatically triggered given the predicted end state (Sec. IV-B). Our approach focuses on generating post-grasp trajectories, relying on pre-generated grasp poses to grasp the object. Subsequently, actions are executed based on a synthesized pose trajectory. For object pose tracking, we reconstruct the object mesh from the training demonstration sequence using BundleSDF [wen2023bundlesdf] and obtain the first frame object mask with SAM6D [lin2024sam] through RGBD observations and the object mesh.
Baselines. We compare with two state-of-the-art 3D-based imitation learning methods: RVT2 [goyal2024rvt] and 3D Diffuser Actor [ke20243d]. Both are language-conditioned, multi-task policies. We use their officially released implementations for retraining on the set of considered tasks. Notably, RVT2 requires multi-camera setup (4 in this case), whereas 3D Diffuser Actor supports single-view, the same as our setup.
Table I: Simulation Results on RLBench. All methods use a single set of weights to perform multiple tasks, conditioned on language description.
@@
Analysis. Our method achieves performance comparable to the multi-view approach RVT2, even with a single camera setting where occlusion often occurs. Additionally, our method excels in high-precision tasks such as Insert Peg and Place Cups, outperforming other baselines. By leveraging pose trajectory representation, our method captures the essential spatial relationship between the task-relevant objects, as well as the intermediate transformations during the process. Furthermore, our method surpasses other baselines in long-horizon tasks like Stack Blocks and Stack Cups, as its object-centric representation naturally decomposes long-horizon, multi-object task learning into a series of sub-policies. However, our method struggles with thin objects (the stick in Drag Stick), small objects occluded by the gripper (the bulbs in Screw Bulb), or a mix of objects with different symmetries (the shape toys in Sort Shape).
V-C Results in Real-world
@@
Baselines. Learning policy from action-less video while adhering to task-specific constraints is challenging. Previous works [wen2023any, vecerik2024robotap, bharadhwaj2024track2act, yuan2024general] exploit point tracking for extracting task-relevant actions from videos. Accordingly, we implement a Point-Tracking baseline, utilizing a heuristic point-tracking method to derive object trajectories from the video. Specifically, we uniformly sample key points from the first frame, track these points using the CoTracker [karaev2023cotracker], apply RANSAC to extract object transformations throughout the sequence, and use these transformations for learning trajectory synthesis. This baseline allows us to compare our approach with existing methods in modeling intermediate states and implicit task constraints.
@@ VI Conclusion
We present SPOT, an object-centric imitation learning method that allows robots to learn from cross-embodiment demonstrations and adhere to task constraints using an intermediate representation of object pose trajectories. Our approach showcases robust real-world manipulation capabilities, outperforming all baselines across various manipulation tasks. Despite its success, SPOT has limitations: 1) The learned policy cannot handle non-prehensile tasks. 2) Our approach relies on 6D pose tracking, which assume object is rigid and may struggle with small, thin, or material-less objects. 3) Object tracking requires reconstructed object mesh, though this has become more accessible with several recent advancements [wen2023bundlesdf, wang2024dust3r].

### v2

@@ IV-B Object Trajectory Diffusion
Our method decouples observation-action pairs in visuomotor policy learning by employing object pose trajectory as an intermediate representation. We model $p(T_{t:t+h}|T_{t})$ to infer a future object pose trajectory $T_{t:t+h}\in SE(3)$ that accomplishes the task, from the current object pose $T_{t}\in SE(3)$.
@@
$A_{t}^{k-1}=\alpha\left(A_{t}^{k}-\gamma\epsilon_{\theta}(O_{t},A_{t}^{k},k)+\mathcal{N}(0,\sigma^{2}I)\right)$ (1)
The process starts with ${A}_{t}^{K}$ sampled from Gaussian noise and is iterated $K$ times. Here, $(\alpha,\gamma,\sigma)$ are the parameters of the denoising scheduler. This process is also called the reverse process and is detailed in previous work [ho2020denoising]. The score function $\epsilon_{\theta}$ is trained by the loss function
@@
In our framework, we model $p(T_{t+1}\mid T_{t})$ as the primary component for policy learning. Rather than predicting future actions, we synthesize the object pose trajectory for task progression. Additionally, our diffusion process is conditioned on the historical pose until now rather than raw sensory input. To adapt this method to our framework, we replace the observation input $O_{t}$ with the current object pose $T_{t}$. Furthermore, to obtain the object pose trajectory, $A_{t}$ is replaced by the transformation $T_{t+1}$. Accordingly, we modified Eq. (1) to
$T_{t}^{k-1}=\alpha\left(T_{t}^{k}-\gamma\epsilon_{\theta}(T_{t},T_{t+1}^{k},k)+\mathcal{N}(0,\sigma^{2}I)\right)$ (3)
@@
Multi-task Language Conditioning. Thanks to the flexibility of diffusion models, we can effortlessly adapt to single-model, multi-task settings by conditioning on other multimodal data, such as language. Specifically, task descriptions are fed into a pretrained CLIP [radford2021learning] model to obtain sentence embeddings, following previous work [shridhar2023perceiver, goyal2023rvt, goyal2024rvt]. These feature embeddings are then concatenated with each object pose as input to the model.
We use DDIM [song2020denoising] as the noise scheduler. The scheduling is set to 100 timesteps during training and 10 timesteps during inference. The number of training epochs is 3000, and the batch size is 128 for all the baselines.
@@ IV-C Closed-loop Policy Execution
To derive the closed-loop manipulation policy $\pi$ from our object-centric representation, we need two key connections: one from observation to object pose, $O_{t}\mapsto T^{obj}_{cam}$, and another from pose trajectory to action plan, $T_{t:t+h}\mapsto A_{t:t+h}$. For the first mapping, we utilize off-the-shelf object pose tracking. For the second mapping, the action plan is derived from the synthesized object trajectory. The complete framework consists of three main components: Object Pose Tracking, which provides real-time object poses at each step; Trajectory synthesis, which takes these poses as input and outputs the future object trajectory for task progression; and finally, Action plan generation, which derives the action plan from the synthesized object trajectory.
Object Pose Tracking. We utilize FoundationPose [wen2024foundationpose], a foundation model for 6D object pose estimation and tracking, which can be applied to novel objects in a zero-shot setting. It tracks all the task-relevant SE(3) object poses throughout the policy rollout to provide the latest feedback to the controller.
In addition to RGBD observations, the tracker requires a 3D object model and 2D detection for the first frame. We obtain object scans using an iPhone equipped with a camera and LiDAR, and object bounding boxes by YOLO-World [cheng2024yolo]. Similarly to demonstration data processing, we convert the estimated source object pose into the target object’s coordinate frame before feeding it into the trajectory diffusion model.
@@
Trajectory synthesizing. At test time, we employ the trained diffusion model (Sec. IV-B) to generate future object trajectories that can accomplish the task. At each step, the model uses the previous (if the observation horizon is greater than 1) and current object poses as inputs to predict future object pose trajectories. We aim to achieve long-horizon planning while remaining reactive to unexpected dynamic uncertainties, such as in-hand object slipping or moving target objects. To this end, we integrate the object pose trajectories synthesized by the diffusion model with receding horizon control [mayne1988receding] to ensure robust action execution. Specifically, the model forecasts future object pose trajectories for $N$ steps, from which we select $K$ steps to execute actions. This design enables our approach to achieve closed-loop control by incorporating updated observations from the object pose tracker.
@@
Action plan generation. The synthesized trajectory indicates the desired object poses to complete the task. To convert the pose trajectory into an executable action plan, we transform the future object poses into end effector subgoals using the transformation $T^{obj}_{EE}\in SE(3)$ from the object pose to the end effector pose. Given the extrinsic camera calibration, we can derive the transformation $T^{obj}_{EE}=T^{EE}_{cam}\cdot T^{cam}_{obj}\in SE(3)$ from the object pose to the end effector pose. Consequently, the action can be derived from the predicted object pose by $A_{t:t+h}=T^{obj}_{EE}\cdot T_{t:t+h}$. The action plan is constructed by following these end-effector subgoals. The policy continues to run until the episode ends and changes the gripper state, as signaled by the end-state prediction.
@@ V Experiments
Our experiments aim to address the following research questions: 1) Is our pose trajectory representation effective for capturing task-related 3D information? 2) Can our object-centric approach learn from action-less data and bridge the embodiment and environmental gap between training and testing? 3) Can our object-centric approach generate trajectories that comply with task-specific motion constraints? 4) Can we democratize demonstration data collection with a minimal hardware setup?
@@ V-A Experimental Setup
Simulation setup. We evaluate baselines on RLBench [james2020rlbench], a standard multi-task manipulation benchmark adopted by previous studies [shridhar2023perceiver, goyal2023rvt, goyal2024rvt, ke20243d]. The task and the robot are simulated using CoppeliaSim [rohmer2013v]. A Franka Panda robot equipped with a parallel jaw gripper is utilized for all tasks. Each task is accompanied by a language description and includes 2 to 60 variations, where the task-relevant objects are in different initial/goal configurations or appearances. We chose 13 manipulation tasks that excluded non-prehensile tasks and tasks with articulated objects, as they are beyond the scope of this work. Following previous work [shridhar2023perceiver, goyal2023rvt], all methods are trained on the same set of 100 demonstrations and tested on 25 unseen task configurations.
Real-world setup. To validate learning from videos in unstructured settings without specific environmental or hardware requirements, we recorded demonstration videos in everyday environments using consumer-grade devices. Each task demonstration was performed by a human without robotics hardware, and filmed in various locations such as the kitchen, living room, or office, ensuring no overlap with the robot’s testing environment. This results in diverse visual backgrounds, lighting conditions, and camera setups that are different from those used during robot testing. The videos were captured with an iPhone equipped with RGB and LiDAR sensors, and all manipulations were performed by human hands. For each task, 8 demonstrations were collected for training, and 10 trials were conducted for each method during evaluation. All real-world experiments were conducted using the same hardware setup: a robot equipped with a Kinova Gen3 7-DoF arm and a forward-facing Intel RealSense D415 RGBD camera mounted at the left shoulder of the robot, which was extrinsically calibrated.
@@ V-B Simulation Results
Implementation details. To collect demonstration trajectories, we use the opening or closing of the gripper as signals to start or terminate a trajectory. Each trajectory corresponds to an episode for single-stage tasks or a stage for multi-stage tasks. For multi-stage tasks, in addition to language embedding, the object trajectory diffusion model is additionally conditioned on a one-hot encoding of the task stage. The next task stage is automatically triggered given the predicted end state (Sec. IV-B). Our approach focuses on generating post-grasp trajectories, relying on pre-generated grasp poses (derived from the demonstration) to grasp the object. Subsequently, actions are executed based on a synthesized pose trajectory. For object pose tracking, we reconstruct the object mesh from the training demonstration sequence using BundleSDF [wen2023bundlesdf] and obtain the first frame object mask with SAM6D [lin2024sam].
Baselines. We compare with two state-of-the-art 3D-based imitation learning methods: RVT2 [goyal2024rvt] and 3D Diffuser Actor [ke20243d]. Both train language-conditioned, multi-task policies. We use their officially-released implementations for retraining on the set of considered tasks. Notably, RVT2 requires a multi-camera setup (4 in this case), whereas 3D Diffuser Actor supports single-view, the same as our setup.
Table I: Simulation Results on RLBench. All methods use a single set of weights to perform multiple tasks, conditioned on language description. Notably, RVT2 uses a four-camera setup, while 3D-DA and Ours use a single-view setup.
@@
Analysis. Our method achieves performance comparable to the multi-view approach RVT2, even in a single-camera setting where occlusion often occurs. Additionally, our method excels in high-precision tasks such as Insert Peg and Place Cups, outperforming other baselines. By leveraging a pose trajectory representation, our method captures the essential spatial relationships between the task-relevant objects, as well as the intermediate transformations during the process. Furthermore, our method surpasses other baselines in long-horizon tasks like Stack Blocks and Stack Cups, as the object-centric representation naturally decomposes long-horizon, multi-object task learning into a series of sub-policies. However, our method struggles with thin objects (the stick in Drag Stick), small objects occluded by the gripper (the bulbs in Screw Bulb), or a mix of objects with different symmetries (the shape toys in Sort Shape).
V-C Real-world Results
@@
Baselines. Learning policies from action-less videos while adhering to task-specific constraints is challenging. Previous works [wen2023any, vecerik2024robotap, bharadhwaj2024track2act, yuan2024general] exploit point tracking for extracting task-relevant actions from videos. Accordingly, we implement a Point-Tracking baseline, utilizing a heuristic point-tracking method to derive object trajectories from the video. Specifically, we uniformly sample key points from the first frame, track these points using CoTracker [karaev2023cotracker], apply RANSAC to extract object transformations throughout the sequence, and use these transformations for learning trajectory synthesis. This baseline allows us to compare our approach with existing methods in modeling intermediate states and implicit task constraints.
@@ VI Conclusion
We present SPOT, an object-centric imitation learning method that allows robots to learn from cross-embodiment demonstrations and adhere to task constraints using an intermediate representation of object pose trajectories. Our approach demonstrates robust manipulation capabilities in both simulation and real-world environments, outperforming all baseline methods across a variety of manipulation tasks.
Limitations. SPOT only considers the SE(3) poses of grasped objects, limiting its ability to handle non-prehensile tasks or dynamic tasks that require regulating the robot’s speed or force. SPOT depends on 6D pose tracking, which assumes the object is rigid. This method may struggle with small, thin, non-rigid, or deformable objects and could be impacted by significant occlusion. Object tracking requires a reconstructed object mesh, though recent advancements have made this more accessible [wen2023bundlesdf, wang2024dust3r].